OpenStudy (anonymous):

log6(x+3)+log6(x+4)=1

OpenStudy (anonymous):

First, we must know the condition of solution. The equation has sense iff \[x+3>0\text{ et } x+4>0\] This is equivalent to : \[x>-3\] We have : \[\log_6(x+3)+\log_6(x+4)=1\iff \log_6\left[(x+3)(x+4)\right]=1 \\\iff (x+3)(x+4)=e\\ \iff x^2+3x+4x+12=e\\ \iff x^2+7x+12-e=0 \] And we have : \[\Delta=49-4\times(12-e)=1+4e\] So : \[x_1=\frac{-7+\sqrt{1+4e}}{2}\simeq -1.78>-3\quad\text{accepted}\\ x_2=\frac{-7-\sqrt{1+4e}}{2}\simeq-5.22<-3 \quad\text{rejected}\]